OpenStudy (anonymous):

You have found the following ages (in years) of all 5 meerkats at your local zoo: 5,2,2,3,14 What is the average age of the meerkats at your zoo? What is the standard deviation? You may round your answers to the nearest tenth. @johnweldon1993 do you know anything about this? :S

OpenStudy (johnweldon1993):

Well first...what is the average?

OpenStudy (anonymous):

Any idea how do calc the avg?

OpenStudy (anonymous):

@johnweldon1993 Looks like you got this?

OpenStudy (anonymous):

5.2?

OpenStudy (johnweldon1993):

yeah no problem @taljaards

OpenStudy (johnweldon1993):

Correct @hemmohgosh5sos so now...for the standard deviation...

OpenStudy (johnweldon1993):

you take the average...(5.2) and you subtract every number up there from it... We want the difference between every number and the average...so for example...with 14 5.2 - 14 = ?

OpenStudy (johnweldon1993):

might want to check your math again @rgalli1

OpenStudy (anonymous):

-8.8

OpenStudy (johnweldon1993):

So we want to answer all of those @hemmohgosh5sos 5.2 - 5 5.2 - 2 5.2 - 2 5.2 - 3 5.2 - 14 answer all those for me :)

OpenStudy (anonymous):

5: 0.2 2: 3.2 2:3.2 3: 2.2 14: -8.8 correct?

OpenStudy (johnweldon1993):

Correct :) so now...we take every one of those results...and square them .2^2 = ? 3.2^2 = ? etc....give me all those results ... :)

OpenStudy (anonymous):

.04 10.24 10.24 4.84 -77.44

OpenStudy (johnweldon1993):

everything but 1 is correct... (-8.2)^2 = 77.44 not negative :) Alright so we have \[\large .04 , 10.24, 10.24, 4.84, 77.44\] Add all of those up...

OpenStudy (anonymous):

oh my bad lol but its 103.16. :) right?

OpenStudy (johnweldon1993):

not...quite what i get...do it 1 more time :) I could be wrong so I will too :)

OpenStudy (johnweldon1993):

Yeah I get 102.8

OpenStudy (johnweldon1993):

is that what you get @hemmohgosh5sos

OpenStudy (anonymous):

i got 103.16 again. hmmm let me try one more time

OpenStudy (johnweldon1993):

You're doing .4 not .04 :P

OpenStudy (anonymous):

you're right, lol my mistake

OpenStudy (johnweldon1993):

:D alright so we have 102.8 we divide that by 5 (yes like taking the average)

OpenStudy (anonymous):

20.56

OpenStudy (johnweldon1993):

right...and this what we have now 20.56 is the variance....to find the standard deviation....we just take the square root of the variance...so \[\large \sqrt{20.56} = ?\]

OpenStudy (anonymous):

4.534

OpenStudy (johnweldon1993):

right! and rounded to the nearest tenth...?

OpenStudy (anonymous):

it would just be 4.5 though right?

OpenStudy (johnweldon1993):

^Indeed :) And that is your answer :)

OpenStudy (anonymous):

thank you again, you're a life saver!!!

OpenStudy (johnweldon1993):

Anytime hun :)

OpenStudy (anonymous):

can i try a few more and you correct me if im wrong? if not its alright!